How far from the y-axis is the center of the curve 2x²+2y²+10x-6y-55=0


1
Expert's answer
2022-01-31T17:54:02-0500

Let us find how far from the "y"-axis is the center of the curve "2x^2+2y^2+10x-6y-55=0."

This equation is equivalent to "x^2+y^2+5x-3y=27.5," and hence to "x^2+5x+(2.5)^2+y^2-3y+(1.5)^2=27.5+(2.5)^2+(1.5)^2."

The last equation is equivalent to "(x+2.5)^2+(y-1.5)^2=36."

Therefore, the point "(-2.5,1.5)" is the center of the circle "2x^2+2y^2+10x-6y-55=0."

It follows that that the distance between the y-axis and the center "(-2.5,1.5)" is equal to "2.5."